OpenStudy (anonymous):

please help! sqrt{xsquared-x-5}=5 A. -5 B. -6 C. 6 D. -5, 6

OpenStudy (anonymous):

\[\sqrt{x^{2}-x-5}=5\] is this your problem ?

OpenStudy (anonymous):

yes!:)

OpenStudy (anonymous):

couldn't find a solution

OpenStudy (anonymous):

Start by squaring both sides.

OpenStudy (anonymous):

i just need the answer please

OpenStudy (anonymous):

Nope, not doing your work for you.

OpenStudy (anonymous):

D. -5,6

OpenStudy (anonymous):

you sure tennis?

OpenStudy (anonymous):

you sure tennis?

OpenStudy (anonymous):

98.25% sure

OpenStudy (anonymous):

So yeah pretty sure :D

OpenStudy (anonymous):

i see where i went wrong teinis is right

OpenStudy (anonymous):

thankyou tennis so much! i have more questions like this would you be willing to help me? :)

OpenStudy (anonymous):

solve for x \[- \sqrt{x}=-1\] A. 0 B. -1 C. 1 D. No solution

OpenStudy (anonymous):

1

OpenStudy (anonymous):

solve for x \[\sqrt{x squared+8x}=4\sqrt{3}\] A. -12 B. -4 C. 4 D. -12, 4

OpenStudy (anonymous):

sorry i did not respond but my little brother pulled my power cord to my computer :( but the answer to that first question is -5,6 because when you square both sides to get rid of that square root you get \[x^2 - x -5 = 5^2=25 \]\[x^2 -x-30=0\]\[(x-6)(x+5)\]so x=-5 or x=6